OpenStudy (countrygirl1431):

Solve the following equation. Then place the correct mixed number in the box provided.

OpenStudy (countrygirl1431):

5x/7=22

OpenStudy (phi):

if this is \[ \frac{5x}{7}= 22 \] multiply both sides by \( \frac{7}{5} \)